OpenStudy (anonymous):

Write the equation for the given graph (-8,8) (8,8)

OpenStudy (anonymous):

at the equation Y=mX+b at "m" is the slope and "b" is the intersection part with Y-axis ;but "slope"= (Y2-Y1)/(X2-X!0 then "m"= (8-8)/(8-(-8))=0 then the line is parallel to Y-axis and its equation has no X-componant ;then the equation is Y=8

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw